# Opinion de-polarization of social networks with GNNs

### v1

@@ Abstract
Nowadays, social media is the ground for political debate and exchange of opinions. There is a significant amount of research that suggests that social media are highly polarized. A phenomenon that is commonly observed is the echo chamber structure, where users are organized in polarized communities and form connections only with similar-minded individuals, limiting them- selves to consume specific content. In this paper we explore a way to decrease the polarization of networks with two echo chambers. Particularly, we observe that if some users adopt a moderate opinion about a topic, the polarization of the net- work decreases. Based on this observation, we propose an efficient algorithm to identify a good set of K users, such that if they adopt a moderate stance around a topic, the polarization is mini- mized. Our algorithm employs a Graph Neural Network and thus it can handle large graphs more effectively than other approaches
@@ 1 Introduction
In the contemporary digital landscape, dominated by platforms like Instagram, Facebook, and TikTok, users find themselves immersed in an ever-expanding sea of content (Bakshy et al., 2015). Despite the sheer volume, the user’s experience usually expands to limited content. The main reason behind this phenomenon lies in the recommendation algorithms employed by these platforms. Those algorithms capture user preferences and interactions and are designed to recommend personalized content. Combined with homophily - the tendency of people to interact with similar-minded individuals. This fact leads to a significant decrease in the variety of posts that appear in the user’s feed.
@@
This narrowing of exposure leads to a broader issue of polarization, where users organize themselves into isolated, ideologically homogeneous communities, commonly referred to as echo chambers. Within these echo chambers, individuals’ pre-existing beliefs are reinforced, often leading to extremism, fanaticism, and a diminished openness to alternative viewpoints. Addressing this challenge has become a central focus in social network analysis, with several studies proposing methods to reduce polarization, some through friend recommendation from outside the echo chamber (Tommasel et al., 2021) while others through recommendation of diverse content (Cinus et al., 2022). However, these approaches do not explicitly account for users’ opinions or quantify the level of polarization in the network. They primarily focus on diversifying content recommendations without a deeper consideration of the underlying opinion dynamics. Our main focus is on networks with two echo chambers, a phenomenon that has been observed in multiple contexts, such as politics (democrats - republicans), and COVID-19 vaccination.
@@
In this work, we adopt the famous opinion evolution model of Friedkin and Johnsen (Friedkin & Johnsen, 1990). According to this model, the opinion that an individual adopts and expresses depends on the individual’s deeply held internal beliefs and the expressed opinions of her social environment. The details of the model are given in Section 3.
@@
This opinion model introduces numerical values as opinions and thus it enables us to quantify the network’s polarization and try to decrease it. Moreover, the change of an opinion triggers a sequence of opinion updates for multiple users in the network, a process that is commonly referred as diffusion. Based on that, the following question rises: Which user’s opinion change, leads to the largest decrease of the total network’s polarization?
The problem that we tackle is connected to this question. Particularly, we want to identify a set of $K$ users such that when their opinion changes (in a proper way), the network’s polarization is minimized.
Finding exactly the opinion value that each of the $K$ users should have, in order to minimize polarization, is a different problem on its own. In this work, we assume that the $K$ users will adopt and express a moderate opinion. As we discuss in Section 3, the zero value opinion represents the moderate opinion for a topic.
In summary, given a limited budget $K$ we want to minimize the polarization of a social network. So we need to use the limited budget $K$ wisely to convince a proper set of users to adopt moderate opinions, such that the polarization is decreased as much as possible. This could be done either through education of users or through curated recommendations that would show to the users the opposite viewpoint about a topic, such as politics, hoping that they would eventually adopt a moderate stance. Alternatively, one might give incentives to users to adopt and express moderate opinions. In all cases the budget $K$ represents that we have limited resources for this purpose.
Answering those questions is challenging, especially in large networks where significant computation is required. Specifically, this is a combinatorial optimization problem which can’t be solved easily. The authors of (Matakos et al., 2017) proposed an algorithm to solve the problem which, although effective, it’s complexity doesn’t allow it to scale to large graphs.
@@ 2 Related Work
A significant body of research has focused on designing interventions aimed at depolarizing opinions in social networks. One common approach involves the use of content-based recommendations that expose users to diverse or opposing viewpoints in an effort to reduce ideological extremity. In (Vendeville et al., 2022) they model the Newsfeed and Wall of every user as a Markov Chain. They suggested adding curated recommendations to the newsfeeds in order to show users more diverse content. The optimal recommendation rate that maximises the average diversity according to their model, can be found by solving a system.
Another approach centers on friend recommendations, where users are encouraged to connect with individuals who hold different opinions. In (Cinus et al., 2022) they proposed a framework to asses the effects of friend-recommendation systems on the echo chamber and polarisation. They found that friend recommendation systems can strengthen echo chambers under certain conditions, while their effect might be negligible depending on the structure of the underlying network. In the context of friends recommendations, one very interesting work is (Tommasel et al., 2021). They use Graph Convolutional Networks (Kipf & Welling, 2017) along with NLP methods to construct representations of users and the echo chambers, which they later use to recommend relevant and diverse friendships.
The most closely related work to ours, is (Matakos et al., 2017), where the authors use the Friedkin and Johnsen opinion model to address the problem of minimizing overall polarization in a network. Specifically, they aim to identify a set of $K$ users whose opinions, when moderated, lead to a reduction in polarization. This paper is motivated and builds significantly on their approach, tackling the same core problem but with a method that is designed to scale more effectively to larger graphs.
@@ 3 Opinion Dynamics and Polarization Metric
We represent a social network as a graph $G(V,E)$, where $V$ is the set of nodes and $E$ is the set of edges. If node $u$ is connected with node $v$ then the tuple $(u,v)\in E$ and $w_{uv}$ is a weight associated with the edge $(u,v)$. Additionally, each node has two features. Namely, there are two opinions related with each node $i$: the internal opinion $s_{i}$ and the expressed (or external) opinion $z_{i}$, based on the Friedkin and Johnsen opinion formation model (Friedkin & Johnsen, 1990). This is a well studied and commonly used opinion model in the literature. According to it, every node $i$ in a social network maintains a fixed internal opinion $s_{i}$, while she updates her external opinion $z_{i}$ based on the following rule:
@@
The expressed opinion $z_{i}$ of node $i$ represents a compromise between the persistent value of $s_{i}$ and the expressed opinions of others to whom $i$ is connected (Bindel et al., 2015). Both $s_{i}$ and $z_{i}$ take values in the range [-1, +1] where +1 implies complete agreement while -1 complete disagreement with a topic (for example COVID19 vaccination). Intuitively, the internal opinion represents strongly held beliefs that come from the individuals education, background, origin and political orientations. On the other hand, the external opinion is the one that the individual adopts and eventually expresses.
It has been shown that if every person $i$ updates her opinion based on Equation 1, then the expressed opinions converge to a unique opinion vector $\mathbf{z}$, where each component $z_{i}$ is the opinion of user $i$ at the equilibrium (Matakos et al., 2017). The steady state vector $\mathbf{z}$ can be computed by running the evolution process until convergence or by solving the following linear system (Bindel et al., 2015):
$\mathbf{(L+I)z=s}\Leftrightarrow\mathbf{z=(L+I)^{-1}s}$ (2)
where we used the fact that $(L+I)$ is a positive definite matrix and thus invertible ($L$ is the Laplacian Matrix and $I$ is the identity matrix).
Previously, we briefly mentioned that $+1$ and $-1$ imply the complete agreement and disagreement to a topic. Other opinion values in the range $[-1,+1]$ are interpreted similarly, based on their sign, while their magnitude is proportional to their extremity. The zero opinion represents a neutral opinion. Thus, it is a natural choice to measure polarization using the polarization index that was introduced at (Matakos et al., 2017) and is defined as:
$\pi(\mathbf{z})=\frac{||\mathbf{z}||_{2}^{2}}{|V|}$ (3)
In the edge case where $\mathbf{z=0}$, all users express neutral opinion and thus the polarization index is minimized: $\pi(\mathbf{z})=0$. We divide by $|V|$ in order to make the index independent of the network’s size.
@@ 4 Network Depolarization
ModerateExpressed: Given a graph $G=(V,E)$, a vector of internal opinions $\mathbf{s}$, the resulting vector of expressed opinions $\mathbf{z}$, and an integer $K$, identify a set $T$ of $K$ nodes such that fixing the expressed opinions of the nodes in $T$ to 0, minimizes the polarization index $\pi$.
This is one of the problems that was introduced in (Matakos et al., 2017), where they constructed a greedy iterative algorithm to solve it. This algorithm is named GreedyExt. In the same paper, it is proved that ModerateExpressed is an NP-hard problem, and subsequently we don’t know if there exists a polynomial algorithm which can solve the problem optimally.
GreedyExt is an iterative algorithm which starts with an empty set $T^{0}$. At each step t the algorithm adds to the existing solution $T^{t-1}$ the node $v$ which when setting $z_{v}=0$ causes the largest decrease $\pi^{t-1}-\pi^{t}$ in the objective function.
In order to find the best node at each timestep the algorithm evaluates every node independently. For each node, we fix her external opinion to zero and compute the new steady state vector. The most straightforward way to do this is to run the evolution process until convergence using the Equation (1). Alternatively, the new steady state vector, after setting and fixing the external opinion of a node $i$, $z_{i}=0$, can be computed through the following process, as explained in (Matakos et al., 2017):
@@
    Update the Laplacian Matrix $\mathbf{L}$ with a row of zeros at the i-th entry
@@
    Use the Equation (2) with the new $\mathbf{s}$ and $\mathbf{L}$.
Both approaches are valid and were shown to be inefficient for large networks. In particular, solving the linear system (2) for one node involves the multiplication of an $n\times n$ matrix with an $n$-dimensional vector, which translates to $O(n^{2})$, where $n=|V|$. This process is repeated for every node and thus the complexity becomes $O(n^{3})$. Repeating $K$ times (because we are looking for $K$ nodes), leads to $O(Kn^{3})$. In (Matakos et al., 2017), they took advantage of linear algebra properties of the opinion model and were able to bring the complexity down to $O(Kn^{2})$, which, although a great improvement, is still inefficient for large graphs.
@@
Algorithm 1 contains the pseudocode for the GreedyExt. Perhaps, one step that needs further explanation is the part where the algorithm isolates the node $v^{*}$ at each timestep. This is done in order to make sure that $z_{v^{*}}=0$, after the algorithm terminates. If we allow for $v^{*}$ to remain connected with other nodes, then in the next iteration of the algorithm, it’s external opinion will depart from zero because the node will actively participate in the opinion evolution process. As a result, only the best node $v^{*}$ of the $K$-th iteration will have $z_{v^{*}}=0$ after the algorithm terminates which is not what we aim for.
@@
5 Algorithms
As we have already mentioned, GreedExt, although effectively decreases polarization, it doesn’t scale to large networks, due to its high computational complexity. Our approach is very similar to GreedyExt. In particular, we try to approximate the computationally heavy parts of the algorithm and achieve similar results in significantly less time. For this purpose we employee a Graph Neural Network (GNN), which is a neural network architecture specifically designed to work with graph data.

### v2

@@ Abstract
Nowadays, social media is the ground for political debate and exchange of opinions. There is a significant amount of research that suggests that social media are highly polarized. A phenomenon that is commonly observed is the echo chamber structure, where users are organized in polarized communities and form connections only with similar-minded individuals, limiting them- selves to consume specific content. In this paper, we explore a way to decrease the polarization of networks with two echo chambers. Existing work based on popular opinion diffusion models in the literature suggests that, convincing a subset of “key” users to adopt a moderate opinion about a topic, can help reduce the polarization of the entire network. Based on this result, we propose an efficient algorithm to identify a good set of K users such that if they adopt a moderate stance around a topic, the polarization is minimized. Unlike existing work that requires analytically solving a large model iteratively, we propose an algorithm based on Graph Neural Networks (GNN) that (i) can significantly reduce the decision time, (ii) achieve close-to-oracle selection accuracy, and (iii) be robust to the choice of opinion diffusion model.
@@ 1 Introduction
In the contemporary digital landscape, dominated by platforms like Instagram, Facebook, and TikTok, users find themselves immersed in an ever-expanding sea of content (Bakshy et al., 2015). Despite the immense volume of information available on almost every topic, the user’s experience usually expands to limited content. This is for a number of reasons. Homophily, the tendency of people to interact with similar-minded individuals, leads users to form friendships (both online and offline) with similar-minded peers. Psychological phenomena such as biased assimilation and selective exposure can further obstruct a user’s exposure to differing opinion content (Greitemeyer et al., 2009) (Frey, 1986). Last but not least, ubiquitous recommendation algorithms that largely guide a user’s attention are designed to recommend content the user will probably agree with, leading to ever increasing polarization on many sensitive topics.
@@
This narrowing of exposure leads to a broader issue of polarization, where users organize themselves into isolated, ideologically homogeneous communities, commonly referred to as echo chambers. Within these echo chambers, individuals’ pre-existing beliefs are reinforced, often leading to extremism, fanaticism, and a diminished openness to alternative viewpoints. Addressing this challenge has become a central focus in social network analysis, with several studies proposing methods to reduce polarization, some through friend recommendation from outside the echo chamber (Tommasel et al., 2021) while others through recommendation of diverse content (Vendeville et al., 2022). Most such works directly try to identify what content or friendship to recommend to a target user. For instance, a framework that predicts the probability of a new friendship given the users’ previous tweets and interactions is proposed in (Tommasel et al., 2021). Such probabilities are used to recommend friendships from opposing echo chambers that are likely to be formed and thus lead to a decrease in the modularity of the network.
@@
Polarization is interwoven with the opinions of the users of a network, which are often represented numerically. In this work, we adopt one of the most widely used opinion evolution models, known as the Friedkin and Johnsen (FK) model (Friedkin & Johnsen, 1990) (Biondi et al., 2023). According to this model, the opinion that an individual adopts and expresses is represented numerically and depends on the individual’s deeply held internal beliefs and the expressed opinions of her social environment. The details of the model are given in Section 3.
@@
Given this interdependence of the opinions of different users that the FK model suggests, the change in one’s opinion triggers a sequence of opinion updates for multiple users in the network, a process that is commonly referred to as diffusion. Based on that, the following question rises: Which user’s opinion change, leads to the largest decrease of the total network’s polarization?
The problem that we tackle is connected to this question. Specifically, the goal of this work is to accurately and efficiently identify a subset of K users such that their adoption of moderate opinion (moderation) minimizes polarization [^1] How this opinion moderation will take place is orthogonal to our work, and could involve education, incentives (e.g. monetary) and other means, as suggested by (Matakos et al., 2017). Without loss of generality, we will assume that “moderating a user” means setting the level of her expressed opinion to 0 which represents the neutral opinion around a topic, as discussed in Section 3. This is a non-trivial task, because multiple factors including the extremity of the user’s opinion and her friendships must be taken into account when examining whether the user is a good candidate to be moderated.
This problem has been considered in the seminal work of (Matakos et al., 2017), where it is proved that it is an NP-hard problem. The same work proposes an iterative greedy algorithm which selects the node whose moderation minimizes polarization at each timestep, and updates the network accordingly. The greedy algorithm is further compared with reasonable heuristics for the problem and outperforms them. The details of this algorithm are presented in Section 5, but its main drawback is its lack of scalability to large networks, as discussed by the authors.
To this end, this work proposes a novel method which utilizes modern Machine Learning methods to enable efficient identification of key nodes to moderate. We propose a similar greedy algorithm which integrates a Graph Neural Network to simplify the most computationally expensive steps, which are essential for selecting the optimal node at each timestep. Compared to the algorithm introduced in (Matakos et al., 2017), our approach constructs the solution set dramatically faster, while incurring only a minimal difference in the final polarization of the network.
The contributions of this work are:
  • We introduce a Graph Neural Network architecture which generates a score that reflects the effectiveness of a user’s moderation in reducing overall polarization, under some opinion model.
  • Our algorithm utilizes this GNN to solve efficiently the NP-hard problem of choosing a subset of K nodes whose moderation minimizes polarization. We demonstrate up to 16x speedup while maintaining 95% of the performance, compared to previous approaches.
  • We evaluate our model on both synthetic and real-world data, demonstrating its robustness in various scenarios.
@@ 2 Related Work
A significant body of research has focused on designing solutions aimed at depolarizing opinions in social networks. One common approach involves the use of content-based recommendations that expose users to diverse or opposing viewpoints in an effort to reduce ideological extremity. In (Vendeville et al., 2022) a theoretical model that describes how content flows through the platform is used. They suggested adding curated recommendations to the users’ newsfeeds in order to show users more diverse content. The optimal recommendation rate of curated content that maximises the average newsfeed’s diversity according to their model, is found via a quadratic program with linear constraints.
Another approach centers on friend recommendations, where users are encouraged to connect with individuals who hold different opinions. A framework to asses the effects of friend-recommendation systems on the echo chamber and polarisation is proposed in (Cinus et al., 2022). They found that friend recommendation systems can strengthen echo chambers under certain conditions, while their effect might be negligible depending on the structure of the underlying network. In the context of friends recommendations, one very interesting work is (Tommasel et al., 2021). They use Graph Convolutional Networks (Kipf & Welling, 2017) along with NLP methods to construct representations of users and the echo chambers, which they later use to recommend relevant and diverse friendships. While our work also utilizes Graph Convolutional Networks, it is different from (Kipf & Welling, 2017) which constructs a friendhsip recommendation system that recommends friends from outside the target user’s echo chamber, and makes no use of an opinion model.
Unlike those works, our approach attempts to tackle the specific problem setup introduced in (Matakos et al., 2017), where under the FK opinion model and a polarization metric, the authors try to select $K$ users, which when moderated lead to the smallest polarization. This paper is motivated and builds significantly on their approach, tackling the same core problem but with a method that is designed to scale more effectively to larger graphs.
@@ 3 Opinion Dynamics and Polarization Metric
We represent a social network as an undirected graph $G(V,E)$, where $V$ is the set of nodes and $E$ is the set of edges. If node $u$ is connected with node $v$ then the tuple $(u,v)\in E$ and $w_{uv}$ is a weight associated with the edge $(u,v)$. Additionally, each node has two features. Namely, there are two opinions related with each node $i$: the internal opinion $s_{i}$ and the expressed (or external) opinion $z_{i}$, based on the Friedkin and Johnsen opinion formation model (Friedkin & Johnsen, 1990). This is a well studied and commonly used opinion model in the literature (Biondi et al., 2023).
According to it, every node $i$ in a social network maintains a fixed internal opinion $s_{i}$ [^2] Fixed internal opinion does not mean that the user doesn’t change her beliefs and only is affected in terms of what she expresses. Instead, it tries to capture some biases and the various psychological factors, background, education, origin etc. that make some users more or less “pliable” to peers’ opinions, while she updates her external opinion $z_{i}$ based on the following rule:
@@
The expressed opinion $z_{i}$ of node $i$ represents a compromise between the persistent value of $s_{i}$ and the expressed opinions of others to whom $i$ is connected (Bindel et al., 2015). Both $s_{i}$ and $z_{i}$ take values in the range [-1, +1] where +1 implies complete agreement while -1 complete disagreement with a topic (for example COVID19 vaccination). The expressed opinion is the actual opinion of the user and according to the FK model it depends both on her “environment” and internal biases that she has.
It has been shown that if every person $i$ updates her opinion based on Equation 1, then the expressed opinions converge to a unique opinion vector $\mathbf{z^{ss}}$, where each component $z_{i}^{ss}$ is the opinion of user $i$ at the equilibrium (Matakos et al., 2017). The steady state vector $\mathbf{z^{ss}}$ can be computed by running the evolution process until convergence or by solving the following linear system (Bindel et al., 2015):
$\mathbf{(L+I)z^{ss}=s}\Leftrightarrow\mathbf{z^{ss}=(L+I)^{-1}s}$ (2)
In the above $L$ is the weighted Laplacian matrix defined as:
$\mathbf{L}=\begin{cases}\sum_{j\in N(i)}w_{ij},&\text{if }i=j\\ -w_{ij}&\text{if }i\neq j\text{ and }(i,j)\in E\\ 0&\text{otherwise}\end{cases}$
and $\mathbf{I}$ is the identity matrix. $\mathbf{L+I}$ is invertible as a positive definite matrix. Notice that based on Equation 2 the steady state vector $\mathbf{z^{ss}}$ depends only on the internal opinions $\mathbf{s}$ and the Laplacian matrix $\mathbf{L}$ of the underlying graph.
Previously, we briefly mentioned that $+1$ and $-1$ imply the complete agreement and disagreement to a topic. Other opinion values in the range $[-1,+1]$ are interpreted similarly, based on their sign, while their magnitude is proportional to their extremity. The zero opinion represents a neutral opinion. Without loss of generality and in order to facilitate comparison between our results and the results of (Matakos et al., 2017), we use the same polarization metric, known as polarization index. Polarization index is defined as:
$\pi(\mathbf{z^{ss}})=\frac{||\mathbf{z^{ss}}||_{2}^{2}}{|V|}$ (3)
In the edge case where $\mathbf{z^{ss}=0}$, all users express neutral opinion and thus the polarization index is minimized: $\pi(\mathbf{z^{ss}})=0$. We divide by $|V|$ in order to make the index independent of the network’s size. Nevertheless, instead of using the $L_{2}$ norm, we could have used other norms of $\mathbf{z}$, such as the $L_{1}$ norm, in order to measure polarization.
@@ 4 Network Depolarization
ModerateExpressed: Given a graph $G=(V,E)$, a vector of internal opinions $\mathbf{s}$, the resulting steady state vector of expressed opinions $\mathbf{z}$, and an integer $K$, identify a set $T$ of $K$ nodes such that fixing the expressed opinions of the nodes in $T$ to 0, minimizes the polarization index $\pi$.
In simple words, this problem requires us to:
@@
    For each node $i\in T$ set $z_{i}=0$
@@
    Compute the new steady state opinion vector $\mathbf{z^{ss}}$, while maintaining $z_{i}=0,\forall i\in T$ (opinions of nodes in $T$ are “anchored” to zero)
  4. 4.
    The lowest $\pi(\mathbf{z^{ss}})$, the better our solution set $T$ is.
Step 3 requires computing the new steady state vector $\mathbf{z^{ss}}$. The most straightforward way to do this is to run the evolution process until convergence using the Equation (1). It is important to manually prohibit nodes in $T$ to depart from zero while running the evolution process.
Alternatively, the new steady state vector, after setting and fixing the external opinions $z_{i}=0,\forall i\in T$, can be computed through the following process, as explained in (Matakos et al., 2017):
  1. 1.
    Set $s_{i}=0,\forall i\in T$
  2. 2.
    Update the Laplacian Matrix $\mathbf{L}$ with a row of zeros at the $i$-th row, $\forall i\in T$
  3. 3.
    Use the Equation (2) with the new $\mathbf{s}$ and $\mathbf{L}$ to get the new $\mathbf{z^{ss}}$.
Next, we will provide a mathematical formulation of the problem ModerateExpressed, but in order to do so, we have to present some notation first:
Let $\mathbf{x}$ be a binary vector (it’s components are either 1 or 0), where $x_{i}=1$ iff node $i\in T$. Furthermore let $\mathbf{D_{x}}$ be the diagonal matrix with the components of $\mathbf{x}$ on it’s main diagonal. Finally $\mathbf{\overline{x}}$ is the complementary vector of $\mathbf{x}$ with components:
$\overline{x_{i}}=\begin{cases}1,\quad\text{if }x_{i}=0\\ 0,\quad\text{otherwise}\end{cases}$
Then we can compute the new $\mathbf{z^{ss}}$ which results from moderating the nodes in $T$ as:
$\mathbf{z^{ss}}=(\mathbf{D_{\overline{x}}}\mathbf{L+I})^{-1}\cdot(\mathbf{s-(s\odot x)})$ (4)
where $\odot$ is the Hadamard product (element-wise multiplication) between $\mathbf{s}$ and $\mathbf{x}$. The matrix $\mathbf{D_{\overline{x}}}$ is used to fill with zeros the rows dictated by the vector $\mathbf{x}$, while the $\mathbf{s-(s\odot x)}$ part of expression (4) is responsible to zero out the corresponding components of $\mathbf{s}$. As we can see $\mathbf{z^{ss}}$ is a function of $\mathbf{x}$ and thus, the polarization index $\pi$ (Eq. 3) is a function of $\mathbf{x}$.
Finally, let:
$f(\mathbf{x})=\frac{||(\mathbf{D_{\overline{x}}}\mathbf{L+I})^{-1}\cdot(\mathbf{s-(s\odot x)})||_{2}^{2}}{|V|}$ (5)
Here $f$ is simply the polarization index $\pi$ of the new steady state vector, expressed as a function of $\mathbf{x}$
With that in mind we are now ready to define formally the problem ModerateExpressed:
$\text{minimize}_{\mathbf{x}}\ f(\mathbf{x})$
subject $\text{subject to}\ \sum_{i=1}^{|V|}x_{i}=K$
With this formulation, we are essentially looking for the binary vector $\mathbf{x}$ with exactly $K$ non zero components, such that the quantity $f\mathbf{(x)}$ in Equation 5 is minimized.
The problem ModerateExpressed is one of the problems that was introduced in (Matakos et al., 2017), and it was proven that it is an NP-hard problem. Subsequently, we don’t know if there exists a polynomial algorithm which can compute an optimal solution. The authors constructed a greedy iterative algorithm which is named GreedyExt in order to solve the problem. Next we present this algorithm.
GreedyExt is an iterative algorithm which starts with an empty set $T^{0}$. At each step $t$ the algorithm adds to the existing solution $T^{t-1}$ the node $v$ which when setting $z_{v}=0$ causes the largest decrease $\pi(\mathbf{z^{ss}_{t-1}})-\pi(\mathbf{z^{ss}_{t}})$ in the polarization index.
In order to find the best node at each timestep the algorithm evaluates every node independently. For each node, we fix her external opinion to zero and compute the new steady state vector along with the decrease in the polarization. After evaluating every node, the one that caused the greatest decrease in the polarization is added to the solution set $T$.
@@
Algorithm 1 contains the pseudocode for the GreedyExt. Perhaps, one step that needs further explanation is the part where the algorithm isolates the node $v^{*}$ at each timestep. This is done in order to make sure that the selected node remains “anchored” to zero throughout the entire algorithm execution. Skipping this step would allow the nodes that have been set to zero in previous timesteps to participate in the evolution process of future timesteps. As a result these nodes would depart from zero which is not what we aim for [^3] In fact this step is required in the case where the new steady state vector $\mathbf{z^{ss}}$ is computed via running the evolution process until convergence, using the Equation (1). In the case where the analytical solution from (2) is used, this step is implicitly taken care of from setting the row of $\mathbf{L}$ for the selected node to zero .
This algorithm, although efficient in finding a good solution, suffers from high computational complexity. Computing the new polarization index after the moderation of each node independently is quite expensive. Specifically, solving the linear system (2) for one node involves the multiplication of an $n\times n$ matrix with an $n$-dimensional vector, which translates to $O(n^{2})$, where $n=|V|$. This process is repeated for every node and thus the complexity becomes $O(n^{3})$. Repeating $K$ times (because we are looking for $K$ nodes), leads to $O(Kn^{3})$.
Fortunately, the authors of (Matakos et al., 2017) observed that for the specific evolution model, $\mathbf{z^{ss}}$ can be computed more efficiently, bringing the complexity of GreedyExt down to $O(Kn^{2})$. Despite a clear improvement in speed, this complexity can still be prohibitively even for moderate values of $K$ and n especially, as we’ll also show in our experiments. Furthermore, variations of the FK model (Biondi et al., 2023) (as well as other opinion evolution models) abound in related work, and there’s no guarantee that such analytical shortcuts (as the one that brought the complexity down to $O(Kn^{2})$) exist for these models.
This work aims to provide a solution to both of these issues. Regarding the complexity of GreedyExt we suggest to incorporate Machine Learning to produce accurate predictions for the decrease of the polarization after moderating a node. With such method, we would be able to replace the computationally expensive part of GreedExt, which has to compute the steady state vector after moderating every node individually. Instead of doing that, with our approach we can select a node at each timestep based on our model’s predictions, significantly speeding up the entire process. Furthermore, we claim that our ML model that produces those predictions is in some way ”opinion-evolution-model-agnostic”. By that, we mean that we could train the ML model on top of different opinion evolution models and predict the corresponding decrease in the polarization. This way we have an efficient version of GreedyExt which is no longer tied to specific opinion-evolution-model’s optimization margins.
@@
5 Proposed Solution
In the end of the previous section we briefly introduced the idea of incorporating a Machine Learning model to enable efficient evaluation of candidate solution nodes. In this section we discuss the details of our suggested approach, which utilizes a Graph Neural Network (GNN). A GNN is a Neural Network architecture that is specifically designed to work with graph data.